OpenStudy (anonymous):

Finding the radius. I haev never understood pi, and would like some explanation. Thanks!

OpenStudy (blurbendy):

Pi is basically this: If you grab a can of soup and some rope, and take the rope and surround the bottom of the can with the rope so that you end up with a circle when you lift the can up, pi is the length around the rope. Make sense?

OpenStudy (anonymous):

Yeah I know that part. I had a picture I want to upload with a question, but for some reason it wouldn't let me even get to my question till now.

OpenStudy (anonymous):

@hba

OpenStudy (blurbendy):

Work out the area of circle that has a radius of 7cm, giving your answer in terms of Pi. All you need to do is substitute r = 7 into A = πr² A = πr² = π × 7² Since 7² = 49 (since 7 × 7 = 49) then you have: = π × 49. = 49π So the final answer is 49π cm² (put the number before Pi)

OpenStudy (blurbendy):

the circumference of a circle is found by multiplying the diameter to π: C = dπ if the radius is 7 yards, then the diameter is 14 yards (diameter is twice the radius) so circumference is C = 14π yards.

OpenStudy (anonymous):

Ok thanks so much! I have never fully understood pi for some reason.

OpenStudy (blurbendy):

No problem. It's a weird concept. Took me a while too.

OpenStudy (anonymous):

:)

OpenStudy (hba):

I am sorry,I guess i am late :/

OpenStudy (anonymous):

:P haha just a wee bit...

OpenStudy (hba):

You can actually use buffon's needle to calculate pi. \[\huge\ \pi \approx\frac{ 2l.n }{ th } \] We found out the best approximation Which is, 355/113. We can also use infinite series .

OpenStudy (hba):

We use, \[\huge\ \pi=\frac{ C }{ d } \] Which is a general way.

OpenStudy (anonymous):

Ok. Well honestly I don't know what some of the stuff is... Thanks though!